Andrew travelled a distance of 39 km by jogging for 4 hours and cycling for 3 hours. He could havetravelledthesamedistancebyjoggingfor7hoursandcyclingfor2hours. Find the speed at which he was jogging and the speed at which he was cycling.
Answer by greenestamps(13200)



4j+3c = 39
7j+2c = 39

Solve the pair of linear equations by any method you choose.

Here is a method you might not have seen that works very nicely for this problem.

Continue the pattern shown by the two equations:
4j+3c = 39
7j+2c = 39 (3 more hours jogging and 1 less hour cycling gives the same total distance)
10j+c = 39 (continue the pattern of 3 more hours jogging and 1 less hour cycling)
13j = 39

So j = 39/13 = 3; his jogging speed is 3km/hr.

Then combining that with the first equation gives c = 9.

ANSWER: jogging 3km/hr; cycling 9km/hr